find the solution of x^2-21=4x. if the solution is a fraction, enter it in simplest form?
Found 2 solutions by checkley75, RAY100:
Answer by checkley75(3666)

x^2-21=4x
x^2-4x-21=0
(x-7)(x+3)=0
x-7=0
x=7 ans.
x+3=0
x=-3 ans.

Answer by RAY100(1637)

x^2-21=4x
subt 4x each side
x^2-4x-21=0
FACTOR
(x+3)(x-7)=0
set (x+3)=0, x=(-3)
set (x-7)=0, x=(7)
check
(-3)^2-21=4(-3)
9-21=-12 ok
(7)^2-21=4(7)
49-21=28 ok